# A Sloppy approach to QSP: XAI enabling fit-for-purpose models

### v2

@@ 3 Discussion
Quantitative Systems Pharmacology (QSP) is a rapidly growing field with significant untapped potential to transform pharmacology[61]. By leveraging detailed mechanistic insights into biological systems, QSP holds promise for accelerating drug development, enabling personalized medicine, and shortening the path from discovery to deployment. Realizing these goals, however, requires state-of-the-art analytical and computational tools that can extract maximum insight from both data and models. In this work, we introduced and validated a QSP workflow that combines literature-derived models with a principled model reduction strategy based on the Manifold Boundary Approximation Method (MBAM). Derbalah et al.[94] list eight characteristics of an “ideal” QSP model reduction system: system agnostic, automatable, mechanistically relevant, accurate, computationally low-cost, extrapolates, requires no [additional] experimental data, and estimates parameters. Traditional MBAM has been limited in QSP deployment by computational expense; however, with advances in computational efficiency, MBAM now fulfills all eight criteria while also offering explainability at all levels of reduction. Our workflow integrates these capabilities into a coherent process, providing a practical and explainable approach to modeling in mainstream QSP applications.
@@
Several key features make our approach particularly well-suited for QSP applications. First, the workflow is grounded in the principles of sloppy modeling, which provide a rigorous theoretical foundation and leverage cutting-edge numerical techniques. Second, our use of a literature-derived model draws on the collective knowledge of the scientific community, helping to reduce bias in model structure and expand coverage of relevant mechanisms. It simultaneously provides a tool to combat Kahneman’s “what you see is all there is” bias[68, 69], and a method for detecting knowledge gaps in the primary literature. Finally, by generating context-specific reduced-order models tailored to target scientific or clinical queries, our approach ensures that predictive accuracy is retained while offering QSP scientists deeper insight into how the system functions.
@@
The algorithmic nature of our workflow makes the modeling process largely automatic, making it a form of eXplainable AI (XAI) for generating fit-for-purpose models. This answers the call from the community to standardized procedures for deciding on model granularity and parameter estimation[63, 62, 60, 6]. As demonstrated, the workflow improves computational efficiency and numerical stability while providing modelers high-level insights into system behavior. By automating key steps in model development, we hope to free QSP scientists from routine technical burdens, allowing them to focus their creative energies on scientific interpretation, hypothesis generation, and experimental design—ultimately helping to realize the full potential of the QSP paradigm.
@@
Taken together, our workflow offers a practical and principled strategy for advancing QSP modeling in complex biological systems. By integrating literature-derived models, sloppy modeling principles, and query-driven model reduction, we provide a framework that is both computationally efficient and scientifically transparent. Importantly, the modular and explainable nature of the workflow makes it well-suited for integration into clinical decision-making pipelines and regulatory review processes, where interpretability, traceability, and predictive reliability are paramount. As QSP continues to expand its role in translational science, tools like this will be essential for scaling insight alongside complexity—and for delivering on the promise of mechanistic modeling in real-world pharmacological applications.
@@ 4.2 Coagulation
where the $k$’s refer to the parameter labels in the original model. The reduced parameters similarlyh have expressions in terms of the original parameters:
@@ 4.3.3 SHIV Calibrated Model
In many QSP models, including those we assimilated into our literature model, some or all parameter values are pinned to literature values. While this is a common strategy for managing parametric complexity, we take the opposite approach in order to stress test our proposed methodology. We obtained best-fit parameters values for each of the three times series of viral load: aggregated untreated, DEWL (responder), and DF06 (nonresponder). We found two statistically equivalent nonresponder fits, DF06A and DF06B, that that use slightly different mechansims and reproduce different qualitative features of the data. DF06A’s simplified drug dynamics capture the early rebound peak, but the long-term increase in V is not monotonic. DF06B’s renormalization of viral clearance leads to smoother long-term growth but underestimates the early peak. In retrospect, DF06B contributed little to further understanding of our model, but including multiple mechanisms is a useful strategy in other contexts, and we keep it here to demonstrate the methodology.
@@ 4.3.4 SHIV Queries and Reduced Models
$\begin{split}\dot{T}&=b_{T}-d_{T}T-f_{I}TV\\ V&=\frac{\widetilde{p_{I}}I}{\widetilde{c}+A}\\ \dot{I}&=f_{I}TV-\widetilde{E}I\\ \dot{\widetilde{E}}&=\frac{\widetilde{b_{E}}I}{K_{B}+I}-\xi\frac{\widetilde{Q}^{n}\widetilde{E}}{1+\widetilde{Q}^{n}}-d_{E}\widetilde{E}\\ \dot{\widetilde{Q}}&=\widetilde{\kappa}I\\ A&=\widetilde{k_{2}}A_{2}\\ T(0)&=\frac{b_{T}}{d_{T}}e^{-\frac{f_{I}}{\widetilde{c}}}\\ I(0)&=\frac{b_{T}}{d_{T}}\left(1-e^{-\frac{f_{I}}{\widetilde{c}}}\right)\\ \widetilde{p_{I}}=\frac{p_{I}}{V_{0}},\widetilde{c}=\frac{c}{V_{0}},\widetilde{k_{2}}=\frac{k_{2}}{KV_{0}},\widetilde{E}&=m_{I}E,\widetilde{b_{E}}=m_{I}b_{E},\widetilde{Q}=\frac{Q}{q_{c}},\widetilde{\kappa}=\frac{\kappa}{q_{c}}\end{split}$ (5)
@@
$\begin{split}\dot{T}&=-\theta_{3}TV\\ V&=\frac{\theta_{1}I}{\theta_{2}+A}\\ \dot{I}&=\theta_{3}TV-EI\\ \dot{E}&=\theta_{4}\\ T(0)&=e^{-\frac{\theta_{3}}{\theta_{2}}}\\ I(0)&=1-e^{-\frac{\theta_{3}}{\theta_{2}}}\\ \theta_{1}=\frac{p_{I}b_{T}}{V_{0}d_{T}},\theta_{2}&=\frac{c}{V_{0}},\theta_{3}=f_{I},\theta_{4}=m_{I}b_{E}\end{split}$ (12)
@@
$d_{Q}\rightarrow 0$ $\widetilde{k_{2}},K\rightarrow\infty$ $KB\rightarrow\infty$ $KB\rightarrow 0$
$\omega_{1}\rightarrow 0$ $k_{1}\rightarrow 0$ $f_{P}\rightarrow 0$
$k_{E}\rightarrow 0$ $m_{I}\rightarrow 1$ $m_{I}\rightarrow 1$
$\widetilde{f}\rightarrow 0$ $k_{2},K\rightarrow\infty$ $V_{0},p_{I},c,k_{2}\rightarrow\infty$
$f_{L}\rightarrow 0$
@@ 4.3.5 SHIV Target Identification
4.3.6 SHIV Uncertainty quantification
QSP models are frequently used to make extrapolative predictions beyond the context of the data from which they are calibrated. For example, a patient model can be queried to investigate the response and side effects of a potential therapeutic intervention. While a parameterized model will always output some prediction, it is crucial to quantify the confidence associated with it in order to properly assess its reliability. While there are many sources of uncertainty in mathematical modeling [96], QSP models are often dominated by parametric uncertainty [97]. Here, we propagate forward the parameter uncertainty resulting from the fitting process to downstream predictions of interest.
A simple method for propagating uncertainty is known as the delta method and is based on asymptotic estimator distributions. Assuming that parameter estimates are close to the Cramér-Rao bound, prediction uncertainty can be calculated using the delta method to construct approximate confidence intervals [93]. The upper and lower bounds are then given by
$g(\hat{\theta})\pm z_{\alpha/2}\ \sqrt{\nabla g\cdot\mathcal{I}^{-1}\cdot\nabla g^{\intercal}}$ (14)
where $g(\theta)$ is the model prediction function being queried, $\mathcal{I}$ is the Fisher information matrix associated with the fitting data, $z_{\alpha/2}$ is the critical value for a given confidence level, and $\hat{\theta}$ are the parameter best-fit estimates. Intuitively, the width of this interval in Eq. (14) is a balance between the parameter uncertainty from the data (quantified by $\mathcal{I}^{-1}$) and the parameter sensitivity of the prediction (quantified by $\nabla g$).
The delta method is based on asymptotic estimates of parameter uncertainty which is a poor approximation for sloppy models. To complement the delta method, we also sample the Bayesian posterior distribution to propagate uncertainty directly to the predictions. The posterior is the conditional probability distribution of the parameters given the observed data, and is calculated using Bayes’ rule:
@@ 4.3.6 SHIV Uncertainty quantification
To sample the Bayesian posterior $\rho(\theta|x)$, we use the affine invariant Markov chain Monte Carlo (MCMC) method [98]. We use an uninformative flat prior centered on the parameter best-fit values. To ensure that the posterior is proper, this flat prior is bounded in each log-parameter at ±$\pm 16$ about the center. Our Markov chain used 48 walkers and ran for 10000 steps after an initial burn-in of 1000 steps. We sampled this chain every 100 steps to eliminate correlations. After the posterior was sampled, uncertainty was propagated by calculating the corresponding model predictions for each sampled parameter set.
We demonstrate our workflow for uncertainty quantification by predicting the effect of bNAb therapy on the patients for which only untreated data is available. Using the model in Eq. 12, we fit the model to the available data and quantify the uncertainty in model queries using both the delta method and the Bayesian method.
5 Acknowledgements

### v3

@@ 3 Discussion
Quantitative Systems Pharmacology (QSP) is a rapidly growing field with significant untapped potential to transform pharmacology[61]. By leveraging detailed mechanistic insights into biological systems, QSP holds promise for accelerating drug development, enabling personalized medicine, and shortening the path from discovery to deployment. Realizing these goals, however, requires state-of-the-art analytical and computational tools that can extract maximum insight from both data and models. In this work, we introduced and validated the Contextualized Reduction for Identifiability and Scientific Precision (CRISP) workflow for QSP that combines literature-derived models with a principled model reduction strategy based on the Manifold Boundary Approximation Method (MBAM). Derbalah et al.[96] list eight characteristics of an “ideal” QSP model reduction system: system agnostic, automatable, mechanistically relevant, accurate, computationally low-cost, extrapolates, requires no [additional] experimental data, and estimates parameters. Traditional MBAM has been limited in QSP deployment by computational expense; however, with advances in computational efficiency, MBAM now fulfills all eight criteria while also offering explainability at all levels of reduction. Our workflow integrates these capabilities into a coherent process, providing a practical and explainable approach to modeling in mainstream QSP applications.
@@
Several key features make our approach particularly well-suited for QSP applications. First, the workflow is grounded in the well-established principles of Information Geometry, which provide a rigorous theoretical foundation and leverage cutting-edge numerical techniques. Second, our use of a literature-derived model draws on the collective knowledge of the scientific community, helping to reduce bias in model structure and expand coverage of relevant mechanisms. It simultaneously provides a tool to combat Kahneman’s “what you see is all there is” bias[68, 69], and a method for detecting knowledge gaps in the primary literature. Finally, by generating context-specific reduced-order models tailored to target scientific or clinical queries, our approach ensures that predictive accuracy is retained while offering QSP scientists deeper insight into how the system functions.
@@
The algorithmic nature of CRISP makes the modeling process largely automatic, yielding a form of eXplainable AI (XAI) for generating fit-for-purpose models. This answers the call from the community to standardized procedures for deciding on model granularity and parameter estimation[63, 62, 60, 6]. As demonstrated, the CRISP workflow improves computational efficiency and numerical stability while providing modelers high-level insights into system behavior. By automating key steps in model development, we hope to free QSP scientists from routine technical burdens, allowing them to focus their creative energies on scientific interpretation, hypothesis generation, and experimental design—ultimately helping to realize the full potential of the QSP paradigm.
@@
Taken together, the CRISP workflow offers a practical and principled strategy for advancing QSP modeling in complex biological systems. By integrating literature-derived models, Information Geometry, and query-driven model reduction, we provide a framework that is both computationally efficient and scientifically transparent. Importantly, the modular and explainable nature of the workflow makes it well-suited for integration into clinical decision-making pipelines and regulatory review processes, where interpretability, traceability, and predictive reliability are paramount. As QSP continues to expand its role in translational science, tools like this will be essential for scaling insight alongside complexity—and for delivering on the promise of mechanistic modeling in real-world pharmacological applications.
@@ 4.2 Coagulation
where the $k$’s refer to the parameter labels in the original model. The reduced parameters similarly have expressions in terms of the original parameters:
@@ 4.3.3 SHIV Calibrated Model
In many QSP models, including those we assimilated into our literature model, some or all parameter values are pinned to literature values. While this is a common strategy for managing parametric complexity, we take the opposite approach in order to stress test our proposed methodology. We obtained best-fit parameters values for each of the three times series of viral load: aggregated untreated, DEWL (responder), and DF06 (nonresponder). We found two statistically equivalent nonresponder fits, DF06A and DF06B, that that use slightly different mechanisms and reproduce different qualitative features of the data. DF06A’s simplified drug dynamics capture the early rebound peak, but the long-term increase in V is not monotonic. DF06B’s renormalization of viral clearance leads to smoother long-term growth but underestimates the early peak. In retrospect, DF06B contributed little to further understanding of our model, but including multiple mechanisms is a useful strategy in other contexts, and we keep it here to demonstrate the methodology.
@@ 4.3.4 SHIV Queries and Reduced Models
$\begin{split}\dot{T}&=b_{T}-d_{T}T-f_{I}TV\\ V&=\frac{\widetilde{p_{I}}I}{\widetilde{c}+A}\\ \dot{I}&=f_{I}TV-\widetilde{E}I\\ \dot{\widetilde{E}}&=\frac{\widetilde{b_{E}}I}{K_{B}+I}-\xi\frac{\widetilde{Q}^{n}\widetilde{E}}{1+\widetilde{Q}^{n}}-d_{E}\widetilde{E}\\ \dot{\widetilde{Q}}&=\widetilde{\kappa}I-d_{Q}\widetilde{Q}\\ A&=\widetilde{k_{2}}A_{2}\\ T(0)&=\frac{b_{T}}{d_{T}}e^{-\frac{f_{I}}{\widetilde{c}}}\\ I(0)&=\frac{b_{T}}{d_{T}}\left(1-e^{-\frac{f_{I}}{\widetilde{c}}}\right)\\ \widetilde{p_{I}}=\frac{p_{I}}{V_{0}},\widetilde{c}=\frac{c}{V_{0}},\widetilde{k_{2}}=\frac{k_{2}}{KV_{0}},\widetilde{E}&=m_{I}E,\widetilde{b_{E}}=m_{I}b_{E},\widetilde{Q}=\frac{Q}{q_{c}},\widetilde{\kappa}=\frac{\kappa}{q_{c}}\end{split}$ (5)
@@
$\begin{split}\dot{\widetilde{T}}&=1-d_{T}\widetilde{T}-f_{I}\widetilde{T}V\\ V&=\frac{\widetilde{p_{I}}\widetilde{I}}{\widetilde{c}+A}\\ \dot{\widetilde{I}}&=f_{I}\widetilde{T}V-\widetilde{E}\widetilde{I}\\ \dot{\widetilde{E}}&=\frac{\widetilde{b_{E}}\widetilde{I}}{\widetilde{K_{B}}+\widetilde{I}}-\xi\frac{\widetilde{Q}^{n}\widetilde{E}}{1+\widetilde{Q}^{n}}-d_{E}\widetilde{E}\\ \dot{\widetilde{Q}}&=\widetilde{\kappa}\widetilde{I}-d_{Q}\widetilde{Q}\\ A&=\widetilde{k_{2}}A_{2}\\ \widetilde{T}(0)&=\frac{1}{d_{T}}e^{-\frac{f_{I}}{\widetilde{c}}}\\ \widetilde{I}(0)&=\frac{1}{d_{T}}\left(1-e^{-\frac{f_{I}}{\widetilde{c}}}\right)\\ \widetilde{p_{I}}=\frac{p_{I}b_{T}}{V_{0}},\widetilde{K_{B}}=\frac{K_{B}}{b_{T}},\widetilde{c}=\frac{c}{V_{0}},\widetilde{k_{2}}&=\frac{k_{2}}{KV_{0}},\widetilde{b_{E}}=m_{I}b_{E},\widetilde{Q}=\frac{Q}{q_{c}},\widetilde{\kappa}=\frac{\kappa b_{T}}{q_{c}},\\ \widetilde{T}=\frac{T}{b_{T}},\widetilde{I}=\frac{I}{b_{T}},\widetilde{E}&=m_{I}E,\end{split}$ (12)
@@
$\omega_{1}\rightarrow 0$ $\widetilde{k_{2}},K\rightarrow\infty$ $KB\rightarrow\infty$ $KB\rightarrow 0$
$k_{E}\rightarrow 0$ $k_{1}\rightarrow 0$ $f_{P}\rightarrow 0$
$\widetilde{f}\rightarrow 0$ $m_{I}\rightarrow 1$ $m_{I}\rightarrow 1$
$f_{L}\rightarrow 0$ $k_{2},K\rightarrow\infty$ $V_{0},p_{I},c,k_{2}\rightarrow\infty$
@@ 4.3.5 SHIV Target Identification
4.3.6 SHIV virtual population generation
QSP models are frequently used to make extrapolative predictions beyond the context of the data from which they are calibrated. For example, a patient model can be queried to investigate the response and side effects of a potential therapeutic intervention. While a parameterized model will always output some prediction, it is crucial to quantify the confidence associated with it in order to properly assess its reliability. While there are many sources of uncertainty in mathematical modeling [98], QSP models are often dominated by parametric uncertainty [99]. Here, we propagate forward the parameter uncertainty resulting from the fitting process to downstream predictions of interest by generating a virtual population.
We generate our virtual population by sampling the Bayesian posterior distribution to propagate uncertainty directly to the predictions. The posterior is the conditional probability distribution of the parameters given the observed data, and is calculated using Bayes’ rule:
@@ 4.3.6 SHIV virtual population generation
To sample the Bayesian posterior $\rho(\theta|x)$, we use the affine invariant Markov chain Monte Carlo (MCMC) method [100]. We use an uninformative flat prior centered on the parameter best-fit values. To ensure that the posterior is proper, this flat prior is bounded in each log-parameter at $\pm 8$ about the center. Our Markov chain used 120 walkers and ran for 10000 steps after an initial burn-in of 1000 steps. We sampled this chain every 100 steps to eliminate correlations. After the posterior was sampled, we calculated the corresponding model predictions for each member of this virtual population.
We demonstrate our workflow for virtual population generation by predicting the effect of bNAb therapy on the patients for which only untreated data is available. Using the model in Eq. 12, we fit the model to the available data and generate a virtual population using the Bayesian method. We then categorize the predictions for each virtual patient according to clinical outcome. Differential HIV infection outcomes have been hypothesized to arise from dynamical bistability [95, 80]. Consistent with this established characterization, we determine the steady state stability of each virtual patient.
For our reduced model, we find that dynamics are exclusively monostable or bistable. We therefore first divide the virtual population by this categorization. Because they have only a single steady state, monostable patients cannot exhibit a long-term response to treatment. Bistable patients, however, can end up in either a high or low viral load steady state. Figure 7a compares the dynamics of a typical monostable patient to a bistable responder. The monostable dynamics show a transient response to treatment, while the responder dynamics demonstrate that treatment has perturbed the system from the unhealthy high steady state to the controlled lower steady state.
Bistable virtual patients can be further divided according to their clinical outcomes both with and without treatment, as shown in Figure 7b-f. Immune patients approach the lower steady state both with and without treatment. Untreated responders reach the high steady state, but switch to the low steady state with treatment. Nonresponders approach the high steady state regardless of treatment. Finally, untreated adverse patients approach the healthy low steady state, but are perturbed by treatment to reach the unhealthy high steady state.
Stability is determined by numerically calculating the system fixed points and counting the number of positive solutions—2 for monostable patients, and 4 for bistable patients. Bistable patient responses are categorized by comparing the viral load at day 1000 to the two stable fixed points and determining which is closer.
5 Acknowledgments